# Modular Periodicity of Random Initialized Recurrences

### v2

@@ 1 Introduction
This paper investigates the period landscapes of both cyclotomic and non-cyclotomic recurrences across three main sections. In section 2, we conjecture explicit counting functions for the number of distinct periods modulo $m$ for various cyclotomic families, including $\Phi_{p}$ (1), $\Phi_{2p}$ (4), $\Phi_{p^{j}}$ (7) and $x^{n}-1$ (9), all supported by extensive computational verification. In section 3, the non-cyclotomic Fibonacci and parity recurrences are explored, revealing unexpected chiral symmetry. Despite having algebraically distinct characteristic roots, these recurrences produce identical period structures for all tested moduli. We classify prime moduli based on their quadratic reciprocity and an additional parameter $\alpha$ (11), observe fractal self-similarity at prime power moduli (15), and conjecture weight preservation under modular extension (17). Furthermore, we establish that both recurrences share an identical minimum distribution governed by Lucas number ratios (20). In section 4, we discuss broader implications and open questions, and suggest that deterministic periodicity in finite rings has potential relations to stochastic growth in random recurrences and to classical number-theoretical objects such as modular forms and L-functions.
@@ 2 Periodicity of Cyclotomic Recurrences
Remark 2.
The conjecture implies that when $p$ divides $m$, there are exactly $p$ fixed points (periods of length 1) and $\frac{m^{p-1}-p}{p}$ periods of length $p$. When $p$ does not divide $m$, there is exactly one fixed point (the zero vector) and $\frac{m^{p-1}-1}{p}$ periods of length $p$.
Example 3.
For $p=5$ and $m=10$ (where $5\mid 10$), we have:
@@
Remark 5.
The conjecture implies that when $p$ and 2 divide $m$, there is one fixed point (the zero vector), $\frac{p-1}{2}$ periods of length 2, $\frac{2^{p-1}-1}{p}$ periods of length p and $\frac{m^{p-1}-2^{p-1}-p+1}{2p}$ periods of length $2p$.
Example 6.
For $2p=10$ and $m=10$ (where $p$ and 2 devide m), we have:
@@
Example 8.
For $p^{j}=3^{2}=9$ and $m=12$ (where $3\mid 12$), we have:
@@ 3 Periodicity of Fibonacci Recurrences
These primes exhibit two periodic lengths: the zero vector and non-trivial periods of length $\pi_{B}(p)$.
@@
All primes satisfying $p\equiv 11,19\pmod{20}$ belong to subclass B2, while those satisfying $p\equiv 1,9\pmod{20}$ can belong to either subclass.
Example 12 (Class A Prime: $p=47$, $\alpha=3$).
The Pisano period is $\pi_{A}(47)=\frac{2(47+1)}{3}=32$. The number of distinct periods is:
$_{A}(47)=\frac{3}{2}(47-1)+1=69+1=70$ (16)
The period structure consists of the zero vector and $69$ periods of length $32$.
Example 13 (Class B1 Prime: $p=89$, $\alpha=2$).
The Pisano period is $\pi_{B}(89)=\frac{89-1}{2}=44$. The number of distinct periods is:
$_{B1}(89)=2\cdot(89+1)+1=180+1=181$ (17)
The period structure consists of the zero vector and $180$ periods of length $44$.
Example 14 (Class B2 Prime: $p=11$, $\alpha=1$).
The Pisano period is $\pi_{B}(11)=\frac{11-1}{1}=10$. The number of distinct periods is:
$_{B2}(11)=1\cdot(11+2)+1=13+1=14$ (18)
The period structure consists of the zero vector, $2$ periods of length $5$ and $11$ periods of length $10$.
@@
The period structure for the Fibonacci recurrence and its parity transform exhibit hierarchical self-similarity between prime and prime power moduli. At each transition $p^{k}\to p^{k+1}$, all existing periods are preserved and each period of length $\ell>1$ at $p^{k}$ generates new periods of length $p\ell$ at $p^{k+1}$, with multiplicities scaling by factor $p$. For class B2 primes, however, the multiplicity for the middle period remains constant at exactly $2\alpha$ across all powers $p^{k}$.
Example 16 (Prime Power Extension for B2 prime $p=19$, $\alpha=1$).
The base structure at $p=19$ is $\{1,2\times 9,19\times 18\}$ and at $p^{2}=361$ it is $\{1,2\times 9,19\times 18,2\times 171,379\times 342\}$. Observe that the middle period length $9$ scales to $171=9\times 19$ while maintaining count $2\alpha=2$, and that the main period length $\pi(19)=18$ scales to $342=18\times 19$ with adapted multiplicity increasing from $19$ to $379$.
@@
Example 18 (Extension from $d=2$ to $m=6$).
The space $F_{2}$ contains periods 0 (weight $\frac{1}{4}$) and $011$ (weight $\frac{3}{4}$). When we extend to $F_{6}$, the period 0 extends to periods 0 and $02240442$ with combined weight $\frac{1}{36}+\frac{8}{36}=\frac{9}{36}=\frac{1}{4}$, while the period $011$ extends to periods $033$ and $011235213415055431453251$ with combined weight $\frac{3}{36}+\frac{24}{36}=\frac{27}{36}=\frac{3}{4}$, preserving the original weights exactly.
Example 19 (Extension from $d=3$ to $m=6$).
The space $F_{3}$ contains periods 0 (weight $\frac{1}{9}$) and $01120221$ (weight $\frac{8}{9}$). Upon extension to $F_{6}$, the period 0 extends to periods 0 and $033$ with combined weight $\frac{1}{36}+\frac{3}{36}=\frac{4}{36}=\frac{1}{9}$, while $01120221$ extends to periods $02240442$ and $011235213415055431453251$ with combined weight $\frac{8}{36}+\frac{24}{36}=\frac{32}{36}=\frac{8}{9}$, again preserving the weight distribution perfectly.
@@
$P(n)=\begin{cases}\frac{1}{4}&\text{if }n=0\\ \frac{1}{\pi}\left(\arctan\left(\frac{L_{n-2}}{L_{n-1}}\right)-\arctan\left(\frac{L_{n}}{L_{n+1}}\right)\right)&\text{if }n>1,\,n\text{ even}\\ \frac{1}{\pi}\left(\arctan\left(\frac{L_{n}}{L_{n+1}}\right)-\arctan\left(\frac{L_{n-2}}{L_{n-1}}\right)\right)&\text{if }n>1,\,n\text{ odd}\end{cases}$ (20)
@@
Remark 21.
The probability $P(n)$ arises from the fact, that two minima (of equal absolute value) only exist if $\frac{a_{0}}{a_{1}}=\frac{L_{n}}{L_{n+1}}$ or $\frac{a_{0}}{a_{1}}=\frac{-L_{n}}{L_{n-1}}$ for Fibonaccis recurrence (or if $\frac{a_{0}}{a_{1}}=\frac{L_{n}}{L_{n-1}}$ or $\frac{a_{0}}{a_{1}}=\frac{-L_{n}}{L_{n+1}}$ for its parity transform). Furthermore, between any two consecutive Lucas ratios (e.g. $\frac{L_{n-2}}{L_{n-1}}$ and $\frac{L_{n}}{L_{n+1}}$) lies exactly one mediant Fibonacci ratio (e.g. $\frac{F_{n-1}}{F_{n}}=\frac{L_{n-2}+L_{n}}{L_{n-1}+L_{n+1}}$), which is the unique initial condition for which one $a_{n}=0$ in the recurrence sequence.
Remark 22 (Connection to Modular Forms and Farey Sequences).
The Lucas bounds and intermediate Fibonacci ratios form matrices in $\mathrm{SL}(2,\mathbb{Z})$, suggesting deep connections to modular forms and the theory of continued fractions. The geometric interpretation via arctangent differences may relate to Farey arcs and the tessellation of the upper half-plane by the modular group.
@@ 4 Perspectives
This paper examined the modular periodicity of randomly initialized Fibonacci and cyclotomic recurrences, revealing previously unknown symmetries. We conjectured explicit counting formulas for cyclotomic recurrences and discovered perfect chiral symmetry between the Fibonacci and parity recurrences. We classified prime moduli based on their quadratic reciprocity, observed fractal self-similarity at prime power moduli, established weight preservation under modular extension and derived probability distributions for sequence minima governed by Lucas ratios.
The framework of random initialization naturally extends beyond our quadratic cases. While we investigated the two Littlewood polynomials $x^{2}\pm x-1$ among quadratics, higher-degree polynomials with restricted coefficients also reveal rich periodicity structures. The most significant examples are the order-6 recurrences $a_{n}=a_{n-3}+a_{n-6}$ and $a_{n}=-a_{n-3}+a_{n-6}$ (which arise as negations from $\Phi_{9}$ and $\Phi_{18}$). Analogous to how the Fibonacci and parity recurrences are negations from $\Phi_{3}$ and $\Phi_{6}$, these order-6 recurrences also exhibit identical periodicity for all tested moduli $m\leq 19$, demonstrating that chiral symmetry might extend beyond the quadratic case. A rigorous proof of such symmetries may require companion matrix theory and conjugation properties under sign transformations. Observations show that classical Fibonacci periods (such as $01120221$ mod 3 and $1342$ and $01123033140443202241$ mod 5) appear within these order-6 recurrences, suggesting hierarchical connections. However, their period structure is more complex: certain primes exhibit five distinct period lengths rather than the two or three observed in the quadratic case, and self-similarity at prime powers follows subtler rules. Whether weight preservation and generalized classification formulas extend to this setting remains an open question. Moreover, both order-6 recurrences can be viewed as three interleaved Fibonacci or parity recurrences respectively (via initial ratios $a_{0}/a_{3}$, $a_{1}/a_{4}$ and $a_{2}/a_{5}$), suggesting that the Lucas ratio framework and connections to modular forms established for the quadratic case may generalize, though deriving explicit minimum distributions for higher-order systems remains unexplored.
Beyond these generalizations, the interplay between different recurrence families appears to encode deep structural information. Preliminary computation reveals systematic patterns in how Fibonacci periods distribute across cyclotomic recurrences: for instance, the period $011$ appears modulo 2 in all $\Phi_{n}$ with $3\mid n$, and modulo 3 in all $\Phi_{3^{j}}$ for $j\geq 2$, while periods from class B2 primes (OEIS A053032, $p\geq 11$) rarely or never appear in cyclotomic landscapes. Whether this reflects deeper structural incompatibility between exponentially growing sequences and unit-root periodicity remains open. Studying complete period landscapes across polynomial families could reveal universal principles governing recurrence periodicity in finite rings. Fekete polynomials, which relate to cyclotomic polynomials and L-functions, as shown in [7], represent a particularly promising candidate for such investigations.
Finally, a central mystery remains: the classification parameter $\alpha$ that determines a prime’s class and governs the exact period count. While we established that the classification is intimately connected to the Legendre symbol $\left(\frac{5}{p}\right)$, with finer structure emerging via congruences modulo 20, the precise value of $\alpha$ for a given prime remains unknown. Understanding whether $\alpha$ relates to higher-order residue properties, Viswanath’s random recurrence through spectral properties, or other arithmetic invariants could eventually lead to an explicit description of the Pisano periods—resolving one of the longstanding open problems in the theory of Fibonacci sequences.

### v3

@@ 1 Introduction
This paper investigates the period landscapes of both cyclotomic and non-cyclotomic recurrences across three main sections. In Section 2, we conjecture explicit counting functions for the number of distinct periods modulo $m$ for various cyclotomic families, including $\Phi_{p}$, $\Phi_{2p}$, $\Phi_{p^{j}}$ and $x^{n}-1$, all supported by extensive computational verification. In Section 3, the non-cyclotomic Fibonacci and parity recurrences are explored, revealing unexpected chiral symmetry. Despite having algebraically distinct characteristic roots, these recurrences produce identical period structures for all tested moduli. We classify their prime moduli via the Legendre symbol and a parameter $\alpha$ governing both period counts and root orders in field extensions, observe fractal self-similarity at prime power moduli and conjecture weight preservation under modular extension. Furthermore, we establish that both recurrences share an identical minimum distribution governed by Lucas number ratios. In Section 4, we discuss broader implications and open questions and suggest that deterministic periodicity in finite rings has potential relations to stochastic growth in random recurrences and to classical number-theoretical objects such as modular forms and L-functions.
@@ 2 Periodicity of Cyclotomic Recurrences
Remark: The conjecture implies that when $p$ divides $m$, there are exactly $p$ fixed points (periods of length 1) and $\frac{m^{p-1}-p}{p}$ periods of length $p$. When $p$ does not divide $m$, there is exactly one fixed point (the zero vector) and $\frac{m^{p-1}-1}{p}$ periods of length $p$.
Example: For $p=5$ and $m=10$ (where $5\mid 10$), we have:
@@
Remark: The conjecture implies, among other things, that when $p$ and 2 divide $m$, there is one fixed point (the zero vector), $\frac{p-1}{2}$ periods of length 2, $\frac{2^{p-1}-1}{p}$ periods of length p and $\frac{m^{p-1}-2^{p-1}-p+1}{2p}$ periods of length $2p$.
Example: For $2p=10$ and $m=10$ (where $p$ and 2 devide m), we have:
@@
Example: For $p^{j}=3^{2}=9$ and $m=12$ (where $3\mid 12$), we have:
@@ 3 Periodicity of Fibonacci Recurrences
These primes exhibit two period lengths: the zero vector and non-trivial periods of length $\pi_{B}(p)$.
@@
All primes satisfying $p\equiv 11,19\pmod{20}$ belong to subclass B2, while those satisfying $p\equiv 1,9\pmod{20}$ can belong to either subclass. The prime $p=5$ is excluded as a degenerate case where the discriminant vanishes modulo $p$.
Remark: The parameter $\alpha$ can be characterized through multiplicative orders in finite field extensions. For class B primes, the roots of $x^{2}-x-1$ lie in $\mathbb{F}_{p}$ with equal orders $(p-1)/\alpha$ (subclass B1) or distinct orders $(p-1)/\alpha$ and $(p-1)/(2\alpha)$ (subclass B2). For class A primes, roots lie in $\mathbb{F}_{p^{2}}$ with order $2(p+1)/\alpha$.
@@
At each transition $p^{k}\to p^{k+1}$ for the Fibonacci and parity recurrences, all period types from $p^{k}$ persist at $p^{k+1}$. Additionally, each period of length $\ell>1$ newly introduced at $p^{k}$ (not present at $p^{k-1}$) generates periods of length $p\ell$ with multiplicity scaled by factor $p$. For class B2 primes with three period lengths, the middle-length periods maintain constant multiplicity $2\alpha$ across all powers.
Example: The base structure (denoted as $\text{multiplicity}\times\text{length}$) for the B2 prime $p=19$ with $\alpha=1$ is $\{1,2\times 9,19\times 18\}$, at $p^{2}=361$ it is $\{1,2\times 9,19\times 18,2\times 171,379\times 342\}$, and at $p^{3}=6859$ it is $\{1,2\times 9,19\times 18,2\times 171,379\times 342,2\times 3249,7219\times 6498\}$.
@@
Example: The space $F_{2}$ contains periods 0 (weight $\frac{1}{4}$) and $011$ (weight $\frac{3}{4}$). When we extend to $F_{6}$, the period 0 extends to periods 0 and $02240442$ with combined weight $\frac{1}{36}+\frac{8}{36}=\frac{9}{36}=\frac{1}{4}$, while the period $011$ extends to periods $033$ and $011235213415055431453251$ with combined weight $\frac{3}{36}+\frac{24}{36}=\frac{27}{36}=\frac{3}{4}$, preserving the original weights exactly. Similarly, the space $F_{3}$ contains periods 0 (weight $\frac{1}{9}$) and $01120221$ (weight $\frac{8}{9}$). Upon extension to $F_{6}$, the period 0 extends to periods 0 and $033$ with combined weight $\frac{1}{36}+\frac{3}{36}=\frac{4}{36}=\frac{1}{9}$, while $01120221$ extends to periods $02240442$ and $011235213415055431453251$ with combined weight $\frac{8}{36}+\frac{24}{36}=\frac{32}{36}=\frac{8}{9}$, again preserving the weight distribution perfectly.
@@
[…]
@@
Remark: The Lucas ratios appearing in the probability formula, together with their mediants—which are Fibonacci ratios of the form $\frac{F_{n-1}}{F_{n}}=\frac{L_{n}+L_{n-2}}{L_{n+1}+L_{n-1}}$—form matrices in $\mathrm{SL}(2,\mathbb{Z})$, suggesting connections to modular forms and continued fraction theory. The geometric interpretation via arctangent differences may relate to Farey arcs and the tessellation of the upper half-plane by the modular group.
@@ 4 Perspectives
This paper examined the modular periodicity of randomly initialized Fibonacci and cyclotomic recurrences, revealing previously unknown symmetries. We conjectured explicit counting formulas for cyclotomic recurrences and discovered perfect chiral symmetry between the Fibonacci and parity recurrences. We classified prime moduli based on the Legendre symbol and a parameter $\alpha$ governing both period counts and multiplicative orders of roots in field extensions. Additionally, we observed fractal self-similarity at prime power moduli, established weight preservation under modular extension, and derived a probability distribution for sequence minima governed by Lucas ratios.
The framework of random initialization naturally extends to higher-degree polynomials. For example, the two order-6 recurrences $a_{n}=\pm a_{n-3}+a_{n-6}$ (negations from $\Phi_{9}$ and $\Phi_{18}$) exhibit chiral symmetry analogous to the Fibonacci case for all tested moduli $m\leq 10$, though their period structure is more complex with up to five distinct lengths for prime moduli. Whether the classification principles established for quadratic recurrences extend to higher-order systems remains an open question.
Preliminary computation reveals systematic embedding of Fibonacci periods into cyclotomic recurrences: class A periods appear in $\Phi_{2(p+1)p^{j}/\alpha}$ modulo $p$, while class B1 periods appear in $\Phi_{(p-1)p^{j}/\alpha}$ modulo $p$. Class B2 primes exhibit more complex behavior, with some showing additional projections to composite moduli. These patterns suggest $\alpha$ encodes structural information about cyclotomic decompositions beyond what the Legendre symbol reveals. Investigating embeddings into other polynomial families — such as Fekete polynomials, which relate to cyclotomic polynomials and L-functions [7] — could reveal universal principles governing recurrence periodicity in finite rings.
Finally, determining $\alpha$ efficiently remains an open problem. While $\alpha$ can be characterized through multiplicative orders of roots in field extensions — requiring factorization of $p\pm 1$ — or through cyclotomic embedding patterns, no efficiently computable form is known. Understanding whether $\alpha$ admits characterization through higher-order residue properties, Galois-theoretic invariants, spectral properties governing random recurrences, or connections to other polynomial families could illuminate the structure of Pisano periods and help resolve one of the longstanding open problems in the theory of Fibonacci sequences.